An ice chest 50 cm long by 40 cm wide by 30 cm high comprised of 3 cm thick Styrofoam having thermal conductivity k = 0.033 W/m°C is filled with ah ice/ water mixture containing 45 kg of ice. The ice / water mixture is separated from the lid by 1 cm of air. The temperature of the surrounding air and the surface on which the ice chest is resting is 30° C, The temperature of the inner surface of the ice chest in contact with the ice / water mixture may be assumed to be 0° C. The heat transfer coefficient at the surface of the ice chest exposed to the air is 20 W/m °C. The thermal contact resistance between the ice chest and the surface on which it is resting is 0.15 m °C/W. Given that the heat of fusion of ice is 333.7 kJ/kg, how long it would take for the ice to melt completely?
Answer #1

Ans. We assume one dimensional steady state heat transfer with constant and unboom Material properties - Conservation of ener

ave I-D the steady state condition top throught and bour sides. So, a = 9 - k.dT À = dx = 0.033 x (8-0)k m . 0.03 m q = 8.8

Under the steady assumption Est- Ein - Eout + Eg = q. Dt. in At - sEst q = 1.50165 X107 ) 5.522 %s = 0.27193951X 1075 & 31.47